Determine sample points
rolling a pair of dice:
a) event of getting a sum of 6
b) even of getting a sum at least 4

Answer by ewatrrr(24785)

Sample space: 36
___1_2_3__4__5__6
1|_2_3_4__5__6__7
2|_3_4_5__6__7__8
3|_4_5_6__7__8__9
4|_5_6_7__8__9_10
5|_6_7_8__9_10_11
6|_7_8_9_10_11_12
...
P(sum of 6) = 5/36
P(sum ≥ 4) = 1 - 3/36 = 11/12